The nominal exchange rate is 80 Indian rupees per dollar. The price of a shirt in India is 1500 rupees. The same shirt sells for $25 in the U.S.

a. What is the real exchange rate? Show your work.

b. Can arbitragers make a profit?

c. If your answer to b is yes, where would they buy and where would they sell? If it is no, why not?
Answer #1

Ans.
Nominal Exchange rate, e = ₹80/$

Price of shirt in India, Pi = ₹1500

Price of shirt in America, Pa = $25

Real Exchange rate, r = e*Pa/Pi

=> r = 80*25/1500 = 1.333

So, dollar is overvalued, so, exchange rate will decrease from the existing ₹80/$ . This means that value of Indian Rupee will increase.

b) Yes, as dollar is overvalued, so, arbitragers can make a profit.

c) As value of dollar is overvalued, so, arbitragers will sell dollar and buy rupee whose value is going to increase in future.

Also, buying shirts from India and selling them in America will make profits for the arbitrager as shirts are cheaper in India and sell expensive in America.

* Please don’t forget to hit the thumbs up button, if you find the answer helpful.

Thank You